# 6DGS: Enhanced Direction-Aware Gaussian Splatting for Volumetric Rendering

### v1

@@ Appendix A Appendix
$D=(d-\mu_{d})^{\top}\Sigma_{d}^{-1}(d-\mu_{d}),$ (9)
and the conditional PDF is given by:
$f_{\text{cond}}=\exp\left(-\lambda_{\text{opa}}\cdot D\right).$ (10)
Table 5 presents the results of our ablation studies on the Synthetic NeRF dataset (Mildenhall et al., 2020). Since this dataset lacks strong view-dependent effects, the influence of $\lambda_{\text{opa}}$ is less significant compared to the 6DGS-PBRT dataset (see Table 4). Compared to our model with $\lambda_{\text{opa}}=0.35$, setting No-$f_{\text{cond}}$ (i.e., $\lambda_{\text{opa}}=0$) slightly degrades image quality and increases the number of Gaussian points. Conversely, when we set No-$\lambda_{\text{opa}}$ (i.e., $\lambda_{\text{opa}}=1$), both the image quality and the number of Gaussians remain comparable. Notably, when we make $\lambda_{\text{opa}}$ a learnable parameter, the model achieves improved image quality with a comparable number of Gaussian points. Furthermore, the No-SH setting results in degraded image quality. Setting the minimum opacity threshold to $\tau=0.005$ (the default value in 3DGS) also degrades the performance and requires more Gaussian points.
Figure 5: More qualitative comparison of methods on the 6DGS-PBRT dataset (zoom in for details).
Figure 5 presents additional qualitative results on the 6DGS-PBRT dataset, illustrating that our 6DGS method consistently achieves superior visual quality compared to 3DGS and N-DG.

### v3

@@ A.3 View-dependent Opacity
$D=(d-\mu_{d})^{\top}\Sigma_{d}^{-1}(d-\mu_{d}),$ (9)
and the conditional PDF is given by:
$f_{\text{cond}}=\exp\left(-\lambda_{\text{opa}}\cdot D\right).$ (10)
A.4 More Ablation Study
@@
Table 8 presents the results of our ablation studies on the Synthetic NeRF dataset (Mildenhall et al., 2020). Since this dataset lacks strong view-dependent effects, the influence of $\lambda_{\text{opa}}$ is less significant compared to the 6DGS-PBR dataset (see Table 4). Compared to our model with $\lambda_{\text{opa}}=0.35$, setting No-$f_{\text{cond}}$ (i.e., $\lambda_{\text{opa}}=0$) slightly degrades image quality and increases the number of Gaussian points. Conversely, when we set No-$\lambda_{\text{opa}}$ (i.e., $\lambda_{\text{opa}}=1$), both the image quality and the number of Gaussians remain comparable. Notably, when we make $\lambda_{\text{opa}}$ a learnable parameter, the model achieves improved image quality with a comparable number of Gaussian points. Furthermore, the No-SH setting results in degraded image quality. Setting the minimum opacity threshold to $\tau=0.005$ (the default value in 3DGS) also degrades the performance and requires more Gaussian points.
A.5 Training Time Comparison
Table 9: Training time (min) comparison of methods on the 6DGS-PBR dataset.
bunny-cloud cloud explosion smoke suzanne ct-scan avg
3DGS 29 38 19 40 41 11 30
N-DG 184 57 56 81 68 131 96
6DGS (Ours) 33 35 31 40 55 20 36
As shown in Table 9, our 6DGS demonstrates competitive training efficiency, requiring only 20% more time than 3DGS (36 vs 30 minutes) while being significantly faster than N-DG (96 minutes). Our training time could be further reduced by implementing Algorithm 1 in CUDA.
A.6 More Qualitative Results
Figure 6 showcases additional qualitative results from the 6DGS-PBR dataset, highlighting the superior visual quality achieved by our 6DGS method compared to 3DGS and N-DG. These visualizations emphasize the ability of 6DGS to capture fine details and complex view-dependent effects while maintaining a high level of geometric fidelity.
Figures 7 and 8 provide qualitative results on three real-world datasets: Deep Blending (Hedman et al., 2021), Tanks & Temples (Knapitsch et al., 2017), and Shiny (Wizadwongsa et al., 2021). Across these datasets, 6DGS demonstrates comparable or superior visual quality to 3DGS. These results further validate the robustness and versatility of 6DGS in handling diverse and challenging scenarios, both synthetic and real-world.
Finally, we provide the Python implementation code for the 6DGS slicing algorithm in Listing 1.
Figure 6: More qualitative comparison of methods on the 6DGS-PBR dataset (zoom in for details).
Figure 7: Qualitative comparison of methods on the Deep Blending (Hedman et al., 2021) and Tanks & Temples (Knapitsch et al., 2017) datasets (zoom in for details).
Figure 8: Qualitative comparison of methods on the Shiny dataset (Wizadwongsa et al., 2021).
1import torch
2
3def slice_6dgs_to_3dgs(L, alpha, mu_p, mu_d, d, iteration, refine_iteratons, lambda_opa=0.35):
4 # Compute the covariance matrix Sigma
5 Sigma = torch.matmul(L, L.T)
6 # Partition Sigma into blocks
7 Sigma_p = Sigma[:, :3, :3]
8 Sigma_pd = Sigma[:, :3, 3:]
9 Sigma_d = Sigma[:, 3:, 3:]
10 # Calculate the conditional mean, covariance, opacity
11 Sigma_d_inv = torch.inverse(Sigma_d)
12 Sigma_regr = torch.matmul(Sigma_pd, Sigma_d_inv)
13 mu_cond = mu_p + torch.matmul(Sigma_regr, x.unsqueeze(-1)).squeeze()
14 Sigma_cond = Sigma_p - torch.matmul(Sigma_regr, Sigma_pd.T)
15 f_cond = torch.exp(-lambda_opa * torch.einsum(’bi,bij,bj->b’, x, Sigma_d_inv, x).unsqueeze(-1))
16 alpha_cond = alpha * f_cond
17
18 s, R = None, None
19 if iteration in refine_iterations:
20 # Perform SVD on the conditional covariance matrix
21 U, S, V = torch.svd(Sigma_cond)
22 # Extract the rotation matrix and scale
23 R = U
24 R[:, 2] *= torch.sign(torch.det(R).unsqueeze(-1))
25 s = torch.sqrt(S)
26 return mu_cond, Sigma_cond, alpha_cond, s, R
Listing 1: Python code for slicing 6DGS to conditional 3DGS